The common stock of KRJ Enterprises is currently selling for $32.78 per share. Last year the company reported total earnings of $341 million and there were 200 million shares of common stock outstanding. Total equity was reported on the balance sheet as $1,990 million. What was the company's price/book ratio? Present your answer rounded to two decimal places (e.g., 20.00.)Please show all steps.
Answer #1

Book value per share = (Total equity)/Number of shares outstanding

Substituting the given values, we get;
= 1990 million/200 million=1990/200=9.95

Price to book ratio=Price per share/Book value per share

=32.78/9.95=3.2944 or 3.29 (Rounded to 2 decimal places)